OpenStudy (anonymous):

Simplify; 28-4r-7r+r^2, i got : 28-3r+r^2.

OpenStudy (kinggeorge):

-4-7=-11 so \[28-4r-7r+r^2= 28 -11r +r^2\]

OpenStudy (anonymous):

thanks<3